First Perfect Square

Given two integers (M and N), write a program to print the first perfect square in a given range.Input


The first line of input will contain a positive integer (M).

The second line of input will contain a positive integer (N).Output


The output should be a single line containing the first perfect square in a given range.Explanation


For example, if the given numbers are M is 4 and N is 16, the perfect squares from 4 to 16 are 4(2 * 2), 9(3 * 3), and 16(4 * 4), as the first perfect square is 4. So the output should be 4.


Sample Input 1

4

16


Sample Output 1

4


Sample Input 2

5

8


Sample Output 2

No Perfect Square


1
Expert's answer
2021-10-31T12:01:49-0400
def ​Squares(a, b):
​
​for i in range (a, b + 1):
​j = 1;
​while j * j <= i:
​if j * j == i:
​yield j*j
​j = j + 1
​i = i + 1

a ​= int(input("Enter lower bound:"))
b ​= int(input("Enter upper bound:"))


x ​= [v for v in Squares(a,b)]
if x:
 ​print(min(x))
ele:
 ​print("No perfect squ